40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 29 14:39:37 2021, Anchal, Summary, BHD, 1Y1 cleared. IY3 ready for C1SUS2 I/O and FE. RemovedStuff.jpeg1Y1.jpeg1Y3.jpeg
    Reply  Tue Nov 2 16:42:00 2021, Paco, Summary, BHD, 1Y1 rack work 
       Reply  Tue Nov 2 18:54:41 2021, Anchal, Paco, Update, BHD, Optical fibres laid for BHD upgrade 
       Reply  Thu Nov 4 15:03:43 2021, Koji, Summary, BHD, 1Y1 rack work 
          Reply  Mon Nov 8 10:13:52 2021, Paco, Summary, BHD, 1Y1 rack work; Sorensens removed 
             Reply  Mon Nov 8 13:13:00 2021, Koji, Summary, BHD, 1Y1 rack work; Sorensens removed 40m_BHD.png
                Reply  Mon Nov 8 15:29:05 2021, Paco, Summary, BHD, 1Y1 rack work; New power for cameras 
                   Reply  Mon Nov 8 18:42:38 2021, Koji, Summary, BHD, Rack Layout / Power Strips rack_plan.pdf
                      Reply  Tue Nov 9 19:02:47 2021, Anchal, Summary, BHD, 1Y0 Populated and 1Y1,1Y0 powered 
Message ID: 16455     Entry time: Mon Nov 8 15:29:05 2021     In reply to: 16454     Reply to this: 16458
Author: Paco 
Type: Summary 
Category: BHD 
Subject: 1Y1 rack work; New power for cameras 

[Paco, Anchal]

In reference to Koji's concern (see previous elog), we have completely removed sorensen power supplies from 1Y1. We added a 12 Volts / 2 Amps AC-to-DC power supply for the cameras and verified it works. We stripped off all unused hardware from shutters and other power lines in the strips, and saved the relays and fuses.

We then mounted SR2, PR3, PR2 Sat Amps, 1Y1 Sat amp adapter, and C1SUS2 AA (2) and AI (3) boards. We made all connections we could make with the cables from the test stand, as well as power connections to an 18 VDC power strip.

ELOG V3.1.3-